Solve for all values of x if x is measured in degrees.
2cos^2x=3sinx
2(1-sin^2x)=3sinx
2-2sin^2x=3sinx
2sin^2x+3sinx-2=0
(2sinx-1)(sinx+2)=0
sinx=1/2 or sinx=-2
Sinx cannot equal -2
so, sinx=1/2
x=[30 +-n(360)]degrees and x=[(150)+-n(360)degrees]
These are all the solutions.
Cheers,
Stan H.
Thanks for your help!

Answer by stanbon(75887)

2cos^2x=3sinx
2(1-sin^2x)=3sinx
2-2sin^2x=3sinx
2sin^2x+3sinx-2=0
(2sinx-1)(sinx+2)=0
sinx=1/2 or sinx=-2
Sinx cannot equal -2
so, sinx=1/2
x=[30 +-n(360)]degrees and x=[(150)+-n(360)degrees]
These are all the solutions.
Cheers,
Stan H.